# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v1

@@ 6.2 WiLDS Datasets
We also test ExPLoRA on the WILDS [29] benchmark, specifically the Camelyon17 [54] and iWildcam [55] datasets, representing domain transfers to medical imagery and wildlife imagery, respectively. Note that the jump from natural images to medical images is larger than for wildlife images, likely because natural image datasets such as ImageNet[56] already contain many images of animals. However, there are little to no medical images such as images of cell tissue in the pre-training datasets for natural image models.
Method PEFT Top 1 Acc.
ConnectLater[57] Full 93.90
@@
D-ExPLoRA-[L]-$r32$ Lin. Probe 94.41
D-ExPLoRA-[L]-$r32$ LoRA-r8 94.21
Table 9: Results on the validation set of Camelyon17
Method PEFT Top 1 Acc.
@@
D-ExPLoRA-[L]-$r32$ Lin. Probe 62.95
D-ExPLoRA-[L]-$r32$ LoRA-r8 68.07
Table 10: Results on the validation set of iWildcam
@@
The iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set of the iWildcam classification task, finding that this outperforms pre-training on the extra-unlabeled set. In table 10, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible loss in knowledge-transfer due to a small domain gap.
@@ 7 Conclusion and Discussion
While effective, there are many aspects of ExPLoRA that deserve further study. The strategy of unfreezing a small number of blocks combines extremely well with PEFT techniques such as LoRA– we hope that future work investigates the reason behind this in further detail. Unresolved questions also include whether other parameter-efficient techniques might be better suited to work with ExPLoRA during pre-training. Further work to evaluate ExPLoRA for natural language domains would be valuable, as would an investigation into whether we can do away entirely with unfreezing a transformer block.
@@
As the scale of models and datasets grows exponentially, access to the computing power necessary to develop and make use of foundation models is increasingly restricted to the hands of a few organizations. Many researchers in academia or smaller companies are then completely dependent on the resources of such organizations in order to leverage ML for their own research or use-cases. Techniques such as PEFT can alleviate this dependence and enable those with much fewer computational resources to perform investigations and customize models for their own use-cases. We hope that ExPLoRA further enables ML practitioners and users to tailor foundation models for their own needs while requiring comparatively few resources. Our hope in doing so is to accelerate the deployment and use of ML for important domains such as sustainability or medicine.
For satellite images, for example, automated analyses that accurately characterize activity on the planet can guide an array of social, economic and environmental policies. Manually curating such observations is time-consuming and expensive, but pre-training foundation models on such data carries its own costs and environmental impact (see appendix C). Thus, we provide a cheaper and more effective way to distill knowledge from foundation models that were already trained on natural images. Insights gained from such a model can further aid researchers and policymakers at a fraction of the cost, enabling more flexible uses of foundation models towards downstream datasets and tasks.
@@ Appendix A Additional Experimental Results
We include further experimental results on downstream datasets as a continuation of section 6.
@@
We consider a land-cover classification dataset used in [23], where each of 244,471 training and 55,529 validation images are paired with one of 66 land cover classes obtained by the USDA’s National Agricultural Imagery Program. In table 11, we first demonstrate similar performance between both natural-image backbones (rows 4 and 5), which surprisingly outperform SatMAE, which is pre-trained on fMoW-RGB. We use ExPLoRA to pre-train from DinoV2 to the training set of this dataset (without labels). Our results (row 6) demonstrate comparable performance, suggesting that for this dataset, domain-specific knowledge may not be highly relevant to successfully solve the task.
@@ Appendix B Training Details
In this section, we describe hyperparameters and hardware configurations used for training our models.
B.1 fMoW RGB
Pre-training
@@
We use the ViT-Large architecture for all experiments. Since raw image sizes vary, the shorter image size is resized to $224$ while preserving aspect ratio, and then a center crop is taken to yield images of size $3\times 224\times 224$, representing the channels, height, and width. For D-ExPLoRA configurations, we use the default setting as used in the code provided by [6] (including for the masking probabilities, relative loss weighting, data augmentations and transforms etc.) Similarly, we use the same settings for M-ExPLoRA pre-training as described in [7, 22].
We use a single NVIDA-RTX 6000 Ada GPU, or 4 NVIDIA-RTX A4000 GPUs for all our experiments, on an academic GPU cluster. For all D-ExPLoRA configurations, we use a batch size of 32, a base learning rate of 1e-3, no weight decay, and a warmup and decaying cosine learning rate scheduler. We train for 200,000 iterations.
@@
For M-ExPLoRA configurations, we use an effective batch size of 1024 (through gradient accumulation), a base learning rate of $41.5e-$ 4, no weight decay, and a warmup and decaying cosine scheduler, with a warmup of 1 epoch, and a total training time of 200 epochs.
Fine-tuning
@@
We use a base learning rate of 1e-3, a cosine scheduler with warmup for 1 epoch, and train for 120 epochs. We use an effective batch size of 256, making use of gradient accumulation if the GPU cannot fit the full batch size in memory. We only use the drop-path augmentation, doing away with mixup and cutmix
Our GPU requirements for fine-tuning are the same as in pre-training.
Linear probing
@@
We adapt the code provided in [6] for linear probing, with a batch size of 256 and a collection of different learning rates: $\left[1e-4,1e-3,5e-3,1e-2,2e-2,5e-2,0.1\right]$. We evaluate both probing on average pooled features as well as on the $\tt{[CLS]}$ token, and also use output features from just the last block, or the last 4 blocks. All numbers reported represent the best validation set accuracy from the best performing configuration.
B.2 fMoW Sentinel
Pre-training
@@
We use the group-channel ViT-L architecture introduced in [22]. We don’t use DinoV2 since there is no such architecture for DinoV2 pre-training. Input images are $13\times 98\times 98$, representing 13 multi-spectral bands. We follow the configuration in [22] of dropping bands B1, B9, B10, and use the same grouping strategy. When loading MAE weights to the ViT-L encoder, the patch embeddings do not match and so the patch embedding and group channel encodings are trained from scratch. All other configuration details are the same as for M-ExPLoRA in section B.1, except that we use a base learning rate of 4.5e-4 for pre-training and train for 50 epochs (given the larger dataset size) on 4 NVIDIA RTX A4000 GPUs for 80 hours.
B.3 Downstream datasets

### v2

@@ 6.2 WiLDS Datasets
We test ExPLoRA on the WILDS (Koh et al., 2021) benchmark, specifically on Camelyon17 (Bandi et al., 2018), iWildcam (Beery et al., 2020) and GlobalWheat David et al. (2020; 2021) datasets, representing domain transfers to medical, wildlife, and agricultural imagery, respectively.
Method PEFT
Top 1
Acc.
CLater [50] Full 93.90
@@
D-[L]-$r32$ Lin. Probe 94.41
D-[L]-$r32$ LoRA-r8 94.21
Table 7: Classification results on the validation set of Camelyon17.
Method PEFT
Top 1
Acc.
@@
D-[L]-$r32$ Lin. Probe 62.95
D-[L]-$r32$ LoRA-r8 68.07
Table 8: Classification results on the validation set of iWildcam.
Method
Top 1
Acc.
AP@
0.5:0.95
AR@
0.5:0.95
ICON [34] 68.9 - -
MAE [26] 82.5 53.8 58.7
DinoV2 [45] 82.3 52.1 57.1
D-[L]-$r64$ 82.7 54.5 59.2
Table 9: Object detection results on the validation set of GlobalWheat. AP and AR stand for average precision and average recall, respectively.
@@
iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set, finding that this outperforms pre-training on the extra-unlabeled set. In table 9, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible loss in knowledge-transfer due to a small domain gap. Likely because natural image datasets $W_{D_{S}}$ such as ImageNet (Deng et al., 2009) used for DinoV2 already contain many images of animals.
GlobalWheat
The GlobalWheat dataset consists of a wheat head object detection task, where each image of a wheat field is associated with bounding boxes on the visible wheat heads David et al. (2020; 2021). ExPLoRA extends pre-training on the training set, and then we run fine-tuning using Detectron2 code for object-detection with ViTs (Wu et al., 2019). ExPLoRA outperforms both fully pre-trained baselines from the WILDS leaderboard and strong VFMs DinoV2 and MAE on top 1 accuracy, average precision, and average recall.
6.3 Analyzing ExPLoRA
Figure 3: The mean of the principal components of the feature map outputted by each ViT block.
Figure 4: Linear probing each patch for position (local information), across all ViT blocks.
Figure 5: Linear probing each patch for classification (global information), across all ViT blocks.
The key design choice of ExPLoRA is to fully train a small subset ${\mathcal{U}}\subset{\mathcal{L}}$ of the ViT, while applying low-rank updates to the remaining frozen layers ${\mathcal{L}}\setminus{\mathcal{U}}$. For parameter-efficiency, we aim to keep $|{\mathcal{U}}|\ll|{\mathcal{L}}|$ and make and informed choice of which layers to unfreeze based on their potential to improve learning during extended pre-training.
We conduct an investigation on 5 models using a sample of ${\mathcal{X}}_{D_{T}}$. These models are DinoV2, D-ExPLoRA-[L]-$r64$, SatMAE, MAE, and M-ExPLoRA-[L]-$r64$. We do the following analyses: (i) PCA to measure the mean and variance of eigenvalues of patch feature vectors for each ViT block, in fig. 3 (ii) linear probing for local or global information (Darcet et al., 2023) by training logistic regression classifiers on each block’s patch feature vectors, to predict either patch position (fig. 4) or image class (fig. 5).
Findings and Unfreezing Strategy for DinoV2
Our analysis reveals that the spectral properties of a block’s feature map (fig. 3) and the ability to retrieve local information from its output patch tokens (fig. 4) are correlated. The classification accuracy for position and the mean of the principal eigenvalues peak in the middle-layers of the model, suggesting that the middle blocks capture local properties of patches (e.g., texture, relative position). Meanwhile, deeper blocks focus on global semantic understanding, as shown by increased classification accuracy for image class prediction in fig. 5. Combined, these results suggest that unfreezing deeper layers, such as ${\mathcal{U}}=\{\texttt{L}\}$, allows the model to better capture global features without overfitting to local details of images of $D_{T}$. This is empirically confirmed in table 3, where linear probing accuracy correlates inversely with the mean eigenvalue of each block (i.e., block 23 > block 22 > block 0 > block 9). The attention maps in fig. 6 further support this, showing that the deeper layers focus more clearly on central objects, while earlier layers (e.g., blocks 9, 10) exhibit more diffuse attention patterns spread around the border.
Findings and Unfreezing Strategy for MAE
For MAE, we see a similar, but less pronounced trend. However, MAE is only trained for reconstruction, and so retains more local information across the ViT’s layers. This is reflected by its lower patch-wise eigenvalues, higher localization accuracy, and lower global accuracies than Dino.
ExPLoRA’s Impact
D-ExPLoRA preserves local information in the middle layers but also improves localization accuracy in the last few layers. Importantly, it also enhances the global information contained in the patches for deeper model layers. This indicates a better understanding of the target domain, as seen in A.2, where ExPLoRA’s attention highlights the central object more clearly.
@@ 7 Conclusion and Discussion
While effective, there are many aspects of ExPLoRA that deserve further study. The strategy of fully training a small amount (or budget) of weights combines extremely well with PEFT techniques such as LoRA– we hope that future work investigates the reason behind this in further detail. Unresolved questions also include whether other parameter-efficient techniques might work better with ExPLoRA during pre-training. Further work to evaluate ExPLoRA for natural language would be valuable, as would an investigation into whether we can do away entirely with unfreezing a transformer block.
@@
As the scale of models and datasets grows exponentially, access to the computing power necessary to develop and make use of foundation models is increasingly restricted to the hands of a few organizations. This leaves many researchers in academia or smaller companies reliant on the resources of such organizations for ML research and applications. Techniques such as PEFT can alleviate this dependence and enable those with fewer computational resources to adapt, investigate, and customize models for their own needs. We hope that ExPLoRA furthers this goal, allowing ML practitioners to tailor foundation models with minimal compute, thus broadening access to powerful ML tools for critical fields like sustainability and medicine.
For example, automated analysis of satellite imagery can inform social, economic, and environmental policies, but manual curation is expensive, and pre-training models on such data has significant costs, both environmental and otherwise (see appendix C). ExPLoRA offers a more efficient way to distill knowledge from existing foundation models trained on natural images, sharply reducing costs while aiding researchers and policymakers and enabling flexible applications in downstream tasks.
@@ Appendix A Additional Experimental Results
We include further experimental results as a continuation of section 6.
A.1 Results on Additional Downstream Datasets
@@
We consider a land-cover classification dataset used in Ayush et al. (2021), where each of 244,471 training and 55,529 validation images are paired with one of 66 land cover classes obtained by the USDA’s National Agricultural Imagery Program. In table 10, we first demonstrate similar performance between both natural-image backbones (rows 4 and 5), which surprisingly outperform SatMAE, which is pre-trained on fMoW-RGB. We use ExPLoRA to pre-train from DinoV2 to the training set of this dataset (without labels). Our results (row 6) demonstrate comparable performance, suggesting that for this dataset, domain-specific knowledge may not be highly relevant to successfully solve the task.
EuroSAT
Method PEFT Top 1 Acc.
SeCo [43] Full 93.14
SatMAE [15] Full 98.98
SatMAE [15] LoRA-r8 98.73
DinoV2 [45] BOFT-b8m2 96.60
M-[1,L]-$r64$ LoRA-r8 98.54
Table 11: EuroSAT validation set results
The dataset contains 27,000 13-band satellite images of 10 classes Helber et al. (2019), sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on the training set of this dataset, and instead use LoRA fine-tuning starting with the pre-trained weights learned in row 8 of table 4. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table 11). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
A.2 Attention Map Visualizations
To aid our analysis in section 6.3, we visualize attention scores for different ViT blocks across multiple models, including DinoV2, D-[L]-$r64$ (i.e. the last row of table 3), the second and third rows of table 3, MAE, SatMAE, and M-[L]-$r64$. These visualizations are shown in fig. 6 for 3 different images from the validation set of fMoW-RGB. Since our models are trained without registers, we truncate attention scores more than 5 standard deviations away from the mean, thus removing artifact attention scores with unusually high values on background patches (Darcet et al., 2023).
Figure 6: Attention maps visualized from the validation set of fMoW-RGB. The models considered, from left to right, are: DinoV2, D-ExPLoRA-[L]-$r64$, Dino with blocks 22,23 unfrozen during extended pretraining, Dino with LoRA-r256 during extended pre-training, MAE, SatMAE, and M-ExPLoRA-[L]-$r64$. We visualize the attention maps at the beginning, middle, and end blocks of the ViT-L.
The visualizations in fig. 6 further support the analysis in section 6.3. For the Dino models, the attention scores of block 9-10 are diffuse and spread around the central object of the image, with quite a few border pixels highlighted. Conversely, the attention scores of the final layers are concentrated more towards the central object. These visualizations further suggest that the middle layers focus on capturing local properties of the images such as texture, while the final layers capture global semantic information such as object-ness. Interestingly, the initial blocks for the Dino models display sparse attention patterns with spikes on seemingly random patches. This might suggest a form of caching to aid the computation of deeper layers that will extract local or global information.
For the MAE models, we see that the original MAE (pre-trained on natural images) seem to highlight more border pixels in the final layers of the ViT. Post extended pre-training with ExPLoRA, the final layers concentrate attention scores on the central object, more closely resembling the patterns of SatMAE (which was fully pre-trained on satellite images). ExPLoRA is thus able to successfully transfer knowledge from its initialized source-domain weights $W_{D_{S}}$ to serve as a foundation model $W_{D_{T}}^{\ast}$ on the new target domain $D_{T}$.
A.3 Convergence and Data Efficiency
Figure 7: Lin. probe accuracy vs. number of training iterations.
Another important question is on ExPLoRA’s data efficiency- i.e. can ExPLoRA achieve good representations on the target domain without requiring many training iterations?
In fig. 7, we plot the linear-probing accuracy against the number of extended pre-training iterations for ExPLoRA (in blue). ExPLoRA improves quickly, requiring between 100-150k extended pre-training iterations to reach optimal performance. As discussed in section 6.1.2, unfreezing additional transformer blocks (in red) fails to achieve the same level of performance while requiring more parameters.
One hypothesis for the effectiveness of pairing unfreezing blocks with LoRA tuning is that the low-rank updates to the ViT backbone “nudge" the sequence of embedded visual tokens from $D_{S}$ to those representing $D_{T}$, which then enables the unfrozen ViT block to efficiently compress global information from the new domain.
A.4 Impact of MAE Decoder Rank
Decoder
Rank $r^{\prime}$
Top 1 Acc.
8 59.75
16 59.77
32 60.15
64 59.21
Table 12: Ablation on M-$\tt{[1,L]}$-$r64$ on the validation set of fMoW-Sentinel. Here, the LoRA rank used for the ViT-L encoder is fixed at $r=64$, while the rank $r^{\prime}$ for MAE decoder is varied.
As outlined in section 5, we initialize the MAE decoder ${\mathcal{L}}_{D}$ with pre-trained weights $W_{D_{S}}$ from He et al. (2022), keeping all decoder weights (except layer norm) frozen during extended pre-training on ${\mathcal{X}}_{D_{T}}$. We apply LoRA with rank $r^{\prime}$ to the $Q,V$ weights of the attention layers in the decoder ${\mathcal{L}}_{D}$, while unfreezing 1-2 blocks ${\mathcal{U}}$ in the ViT encoder ${\mathcal{L}}$ and applying LoRA with rank $r$ to the remaining layers ${\mathcal{L}}\setminus{\mathcal{U}}$ (algorithm 1).
We evaluate ExPLoRA with M-$\tt{[1,L]}$-$r64$ on fMoW-Sentinel, using a fixed encoder LoRA rank $r=64$, unfreezing blocks ${\mathcal{U}}=\{\tt{1,L}\}$, and varying the decoder rank $r^{\prime}$ We then fine-tune the resulting model with LoRA $r=8$ and measure the highest top 1 accuracy on the validation set of fMoW-Sentinel. Results in table 12 show that increasing $r^{\prime}$ up to 32 improves fine-tuning performance, but performance declines when $r^{\prime}=64$, dropping $\downarrow 0.94\%$. This suggests that balancing the unfrozen parameters between the ViT encoder ${\mathcal{L}}$ (used for fine-tuning) and the MAE decoder ${\mathcal{L}}_{D}$ (discarded post pre-training) is crucial. Larger $r^{\prime}$ may improve the decoder’s ability without benefiting the learned representations of ${\mathcal{L}}$. This issue doesn’t arise in DinoV2, as the Dino-iBOT shared head is fully trained since it isn’t provided by Oquab et al. (2023).
A.5 Impact of ViT backbone size
Method Arch.
Top 1 Acc.
Last 1/Last 4
DinoV2 [45] ViT-B 63.62/65.90
DinoV2 [45] ViT-L 67.60/69.00
DinoV2 [45] ViT-G 70.07/70.36
D-[12]-$r64$ ViT-B 74.72/75.11
D-[24]-$r64$ ViT-L 76.86/77.48
D-[32]-$r32$ ViT-G 77.29/77.79
Table 13: Linear probing results on the validation set fMoW-RGB, where we vary the size of the ViT encoder ${\mathcal{L}}$ from ViT-B, ViT-L, and ViT-G. “Last 1/Last 4” refers to using the output representation from just the last 1 or the last 4 ViT layers, respectively.
We also test the impact of the ViT backbone for ExPLoRA, varying the architecture for DinoV2 from ViT-B (86M, $\tt{L}=12$ layers, embedding dimension 768), ViT-L (303M parameters, $\tt{L}=24$ layers, embedding dimension 1024), and ViT-G (1100M parameters, $\tt{L}=40$ layers, embedding dimension 1280) for extended pre-training on fMoW-RGB. The ExPLoRA models we compare against are D-[12]-$r64$ for ViT-B, D-[24]-$r64$ for ViT-L, and D-[32]-$r32$ for ViT-G. We unfreeze the 12th, 24th, and 32nd layers for each of ViT-B, ViT-L, and ViT-G, picking these layers by extending the analysis from section 6.3 to ViT-B and ViT-G. We find that the 12th (last layer) for ViT-B and the 32nd (out of 40) layer for ViT-G output representations with low mean eigenvalues compared to other layers, thus presenting good candidates for unfreezing.
In table 13, we see that as expected, ViT-G performs the best, but is only $\uparrow 0.31\%$ better in top 1 accuracy compared to ViT-L, while using many more parameters. On the other hand, we see the highest impact for ExPLoRA on ViT-B, where the top 1 accuracy improves by $\uparrow 9.21\%$ over the original DinoV2 ViT-B. These results further demonstrate the effectiveness and efficiency of ExPLoRA as a powerful technique to create unsupervised foundation models for new visual domains.
@@ Appendix B Training Details
In this section, we describe hyperparameters and hardware configurations used for our models.
B.1 Pre-Training
@@
We use the ViT-Large architecture for all experiments. Since raw image sizes vary, the shorter image size is resized to $224$ while preserving aspect ratio, and then a center crop is taken to yield images of size $3\times 224\times 224$, representing the channels, height, and width. All pre-training is done on a single NVIDA-RTX 6000 Ada GPU, or 4 NVIDIA-RTX A4000 GPUs on an academic GPU cluster.
ExPLoRA with DinoV2
Most of the hyperparameters for D-ExPLoRA follow the defaults set by Oquab et al. (2023). That is, local (small) crops are between 5%-32% of the original image and are resized to 98x98 pixels, and global (large) crops are greater than 32% of the image and resized to 224x224 pixels. We share the parameters of the Dino-iBOT linear head (3 layers), with a bottleneck dimension of 256, a hidden dimension of 2048, and an output dimension of 65536, initialized from scratch. For Dino, we use Sinkhorn-Knopp (Caron et al., 2020) centering and Koleo (Delattre & Fournier, 2017) regularization with a weight of 0.1. For iBOT, we use masking ratios between 0.1 and 0.5 to mask half of the samples in the batch. The teacher model uses an initial EMA rate of 0.994, with a cosine warmup to 1.000 by the end of training. The teacher warmup and final temperatures are 0.04 and 0.07. The linear Dino-iBOT head is frozen for the first 3k training iterations. We train with the AdamW optimizer (no weight decay), with a base learning rate of $2\times 10^{-3}$ that is varied with a linear warmup and cosine decay schedule. Training is completed within 200,000 iterations, with a batch size of 32 and with 32 gradient accumulation steps (equalling an effective batch size of 1024), and with an epoch length set to 1000.
ExPLoRA with MAE
@@
Most of the hyperparameters we use for M-ExPLoRA pre-training follow those in He et al. (2022); Cong et al. (2022). We use an effective batch size of 1024 (through gradient accumulation), a base learning rate of $4.5\times 10^{-4}$, no weight decay, and a warmup and decaying cosine scheduler, with a warmup of 1 epoch, and a total training time of 200 epochs. We use a masking ratio of $0.75$ and we use the norm_pix_loss flag for the MSE loss.
B.2 PEFT Fine-Tuning
@@
We fine-tune using 4 NVIDIA-RTX A4000 GPUs. We use a base learning rate of $10^{-3}$, a cosine scheduler with warmup for 1 epoch, and train for 120 epochs. We use an effective batch size of 256, making use of gradient accumulation if the GPU cannot fit the full batch size in memory.
For data augmentations, we only use the drop-path augmentation (Larsson et al., 2016) at a rate of 0.2, with no dropout, mixup, or cutmix. We note that the original LoRA configuration outperforms other PEFT techniques when paired with the drop-path regularization technique. For example, we find that BOFT does not pair well with drop-path, instead performing most effectively with a custom multiplicative dropout technique (Liu et al., 2023). We include the result with the best hyperparameter configuration for each row in table 1.
B.3 Linear Probing
@@
We use a single NVIDIA-RTX A4000 GPU for linear probing. We adapt the code provided by Oquab et al. (2023) for linear probing, with a batch size of 256 and a collection of different learning rates: $\left[1\times 10^{-4},1\times 10^{-3},5\times 10^{-3},1\times 10^{-2},2\times 10^{-2},5\times 10^{-2},1\times 10^{-1}\right]$. We evaluate both probing on average pooled features as well as on the $\tt{[CLS]}$ token, and also use output features from just the last block, or the last 4 blocks. All numbers reported represent the best validation set accuracy from the best performing configuration.
B.4 Multi-Spectral Images
@@
We use the group-channel ViT-L architecture introduced in Cong et al. (2022). We don’t use DinoV2 since there is no such architecture for DinoV2 pre-training. Input images are $13\times 98\times 98$, representing 13 multi-spectral bands. We follow the configuration in Cong et al. (2022) of dropping bands B1, B9, B10, and use the same grouping strategy. When loading MAE weights to the ViT-L encoder, the patch embeddings do not match and so the patch embedding and group channel encodings are trained from scratch. All other configuration details are the same as for M-ExPLoRA in section B.1, except that we use a base learning rate of $4.5\times 10^{-4}$ for pre-training and train for 50 epochs (given the larger dataset size) on 4 NVIDIA RTX A4000 GPUs for 80 hours.
Fine-tuning details are the same as in B.2.
B.5 Downstream datasets